1. When 9.0 g of hydrogen gas, H2, reacts with oxygen gas, O2, 73.0 g of water is produced.    


2H2 + O2 --> 2H2O 


a) What is the theoretical yield of water? 

b) What is the percentage yield?


2. Powdered zinc metal reacts with sulphur (S8) when heated to produce zinc sulphide.


8Zn + S8 --> 8ZnS


a) What mass of product can be produced with 25 g of zinc?

b) What is the actual yield if the percentage yield was 80%?


3. The thermite reaction has been used to weld railroad rails, make bombs, and ignite solid rocket fuel. The equation of the reaction is:


Fe2O3 + 2Al --> 2Fe + Al2O3


a) What is the theoretical yield of iron if 100.0 g of Fe2O3 is used?

b) What is the percentage yield if 52.3 g of iron is actually produced?


1
Expert's answer
2022-01-26T03:46:03-0500

1. When 9.0 g of hydrogen gas, H2, reacts with oxygen gas, O2, 73.0 g of water is produced.    

2H2 + O2 → 2H2

a) What is the theoretical yield of water? 

b) What is the percentage yield?


Solution (1):

Calculate the moles of hydrogen gas (H2):

The molar mass of H2 is 2.016 g/mol

Hence,

Moles of H2 = (9.0 g H2) × (1 mol H2 / 2.016 g H2) = 4.4643 mol H2


Balanced chemical equation:

2H2 + O2 → 2H2O

According to stoichiometry:

2 mol of H2 produces 2 mol of H2O

Thus, 4.4643 mol of H2 produces:

(4.4643 mol H2) × (2 mol H2O / 2 mol H2) = 4.4643 mol H2O


Calculate the theoretical yield of water (H2O):

The molar mass of H2O is 18.0153 g/mol

Hence,

Mass of H2O = (4.4643 mol H2O) × (18.0153 g H2O / 1 mol H2O) = 80.426 g H2O = 80.43 g H2O

The theoretical yield of water (H2O) is 80.43 grams


Calculate the percentage yield:

%yield = (actual yield / theoretical yield) × 100%

%yield = (73.0 g / 80.43 g) × 100% = 90.76%

The percentage yield is 90.76%


Answer (1):

a) The theoretical yield of water is 80.43 grams

b) The percentage yield is 90.76%




2. Powdered zinc metal reacts with sulphur (S8) when heated to produce zinc sulphide.

8Zn + S8 → 8ZnS

a) What mass of product can be produced with 25 g of zinc?

b) What is the actual yield if the percentage yield was 80%?


Solution (2):

Calculate the moles of zinc (Zn):

The molar mass of Zn is 65.38 g/mol

Hence,

Moles of Zn = (25 g Zn) × (1 mol Zn / 65.38 g Zn) = 0.3824 mol Zn


Balanced chemical equation:

8Zn + S8 → 8ZnS

According to stoichiometry:

8 mol of Zn produces 8 mol of ZnS

Thus, 0.3824 mol of Zn produces:

(0.3824 mol Zn) × (8 mol ZnS / 8 mol Zn) = 0.3824 mol ZnS


Calculate the mass of product (ZnS):

The molar mass of ZnS is 97.474 g/mol

Hence,

Mass of ZnS = (0.3824 mol ZnS) × (97.474 g ZnS / 1 mol ZnS) = 37.27 g ZnS

The theoretical mass of product (ZnS) is 37.27 grams


Calculate the actual yield:

%yield of product = 80%

%yield = (actual yield / theoretical yield) × 100%

Hence,

actual yield of product (ZnS) = theoretical yield × (%yield / 100%)

actual yield of product (ZnS) = (37.27 g) × (80% / 100%) = 29.82 g

The actual yield of product (ZnS) is 29.82 grams


Answer (2):

a) The mass of product is 37.27 grams

b) The actual yield of product is 29.82 grams




3. The thermite reaction has been used to weld railroad rails, make bombs, and ignite solid rocket fuel. The equation of the reaction is:

Fe2O3 + 2Al → 2Fe + Al2O3

a) What is the theoretical yield of iron if 100.0 g of Fe2O3 is used?

b) What is the percentage yield if 52.3 g of iron is actually produced?


Solution (3):

Calculate the moles of Fe2O3:

The molar mass of Fe2O3 is 159.69 g/mol

Hence,

Moles of Fe2O3 = (100.0 g Fe2O3) × (1 mol Fe2O3 / 159.69 g Fe2O3) = 0.6262 mol Fe2O3


Balanced chemical equation:

Fe2O3 + 2Al → 2Fe + Al2O3

According to stoichiometry:

1 mol of Fe2O3 produces 2 mol of Fe

Thus, 0.6262 mol of Fe2O3 produces:

(0.6262 mol Fe2O3) × (2 mol Fe / 1 mol Fe2O3) = 1.2524 mol Fe


Calculate the theoretical yield of iron (Fe):

The molar mass of Fe is 55.845 g/mol

Hence,

Mass of Fe = (1.2524 mol Fe) × (55.845 g Fe / 1 mol Fe) = 69.94 g Fe

The theoretical yield of iron (Fe) is 69.94 grams


Calculate the percentage yield:

%yield = (actual yield / theoretical yield) × 100%

%yield = (52.3 g / 69.94 g) × 100% = 74.78%

The percentage yield is 74.78%


Answer (3):

a) The theoretical yield of iron is 69.94 grams

b) The percentage yield is 74.78%